OpenStudy (anonymous):

MEDAL + FAN Which fraction is greater than -7/15 A: -5/9 B: -6/11 C: -1/2 D -4/9

OpenStudy (kendricklamar2014):

-7/15 = -0.466667 A. -5/9 = -0.555556 B. -6/11 = -0.545455 C. -1/2 = -0.5 D. -4/9 = -0.444444 Can you tell me which one is greater than -0.46 @Sabu

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

?

OpenStudy (kendricklamar2014):

No, for negative numbers the smaller numbers are gretaer than the larger numbers. Example: -4 is greater than -6

OpenStudy (anonymous):

oh so D? @KendrickLamar2014

OpenStudy (kendricklamar2014):

You got it!

OpenStudy (anonymous):

Thanks Man

OpenStudy (kendricklamar2014):

No Problem